Hockey teams recieve 2 points when they win and 1 point when they tie. One season, a team won a championship with 48 points. they won 9 more games than they tied.
How many wins and how many ties did the team have?

Answer by Nate(3500)

games won = w
games tied = t
w = t + 9
2w + t = 48
Substitute: w = t + 9
2w + t = 48
2( t + 9 ) + t = 48
2t + 18 + t = 48
3t = 30
t = 10
Substitute: t = 10
w = t + 9
w = 10 + 9
w = 19
They won 19 games and tied 10 times.